Ten years from now Amy will be three times as old as she was four years ago.
I got x+10=3x-4. After I solved I got x=7 So her age would be 7+10=17 years right?

Answer by TimothyLamb(4379)

a + 10 = 3(a - 4)
a + 10 = 3a - 12
22 = 2a
---
a = 11
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php